\begin{thebibliography}{s2}
 \addtolength{\itemsep}{-1em}
\bibitem{ash} M.S. Ashbaugh,
Isoperimetric and universal inequalities for eigenvalues, in
Spectral theory and geometry (Edinburgh, 1998) E. B. Davies, and Yu
Safarov eds., London Math. Soc. Lecture Notes vol. 273, Cambridge
Univ. Press, Cambridge, 1999. pp. 95-139.

\bibitem{ash04} M.S. Ashbaugh,
On universal inequalities for the low eigenvalues of the buckling
problem. (English summary) Partial differential equations and
inverse problems, 13-31, Contemp. Math., 362, Amer. Math. Soc.,
Providence, RI, 2004.

\bibitem{cheng4} Q.M. Cheng, H.C. Yang, Inequalities for eigenvalues
of a clamped plate problem, Trans. Amer. Math. Soc. 358 (2006)
2625-2635.

\bibitem{chengyang06} Q.M. Cheng, H.C. Yang,
Universal bounds for eigenvalues of a buckling problem, Commun.
Math. Phys. 262 (2006) 663-675.

\bibitem{harrell} A. EI Soufi, E.M. Harrell and S. Ilias,
Universal inequalities for the eigenvalues of Laplace and
Schr\"{o}dinger operators on submanifolds, Trans. Amer. Math. Soc.
 361 (2009)  2337-2350.

\bibitem{harrell97} E.M. Harrell, J. Stubble,
On trace identities and universal eigenvalue estimates for some
partial differential operators, Trans. Amer. Math. Soc.  349 (1997)
1797-1809.


\bibitem{hile1} G.N. Hile, R.Z. Yeh, Inequalities for eigenvalues
of the biharmonic operator,  Pacific J. Math. 112 (1984) 115-133.

\bibitem{huang09}
G.Y. Huang, X.X. Li and R.W. Xu, Extrinsic estimates for the
eigenvalues of Schr\"{o}dinger operator, to appear in Geom.
Dedicata.

\bibitem{huangchen}
G.Y. Huang, W.Y. Chen, Inequalities of eigenvalues for the bi-kohn
Laplacian on the Heisenberg group, to appear in  Acta Math. Sci.
Ser. B Engl. Ed. 2009, vol 29(6).

\bibitem{levitin}
M. Levitin, L. Parnovski, Commutators, spectral trace identities,
and universal estimates for eigenvalues, J. Funct. Anal. 192 (2002)
425-445.


\bibitem{ppw2} L.E. Payne, G. P\'{o}lya and
H.F. Weinberger, On the ratio of consecutive eigenvalues, J. Math.
Phys. 35 (1956), 289-298.

\bibitem{wangxia07}
Q.L. Wang, C.Y. Xia, Universal inequalities for eigenvalues of the
buckling problem on spherical domains, Commun. Math. Phys. 270
(2007) 759-775.

\bibitem{wang07} Q.L. Wang, C.Y. Xia, Universal bounds for
eigenvalues of the biharmonic operator on Riemannian manifolds, J.
Funct. Anal. 245 (2007) 334-352.



\end{thebibliography}

\documentclass[11pt,twoside]{article}
\usepackage{latexsym,amsmath,amsthm,amssymb,amsfonts}
\usepackage{mathrsfs}
\usepackage{amscd}
\usepackage{cite}

\usepackage[ansinew]{inputenc}
\numberwithin{equation}{section}

\newtheorem{theorem}{\bf Theorem}[section]

\newtheorem{remark}[theorem]{\bf Remark}
\newtheorem{lemma}[theorem]{\bf Lemma}

\newtheorem{corollary}[theorem]{\bf Corollary}
\catcode`@=11

\newskip\plaincentering \plaincentering=0pt plus 1000pt minus 1000pt
\def\@plainlign{\tabskip=0pt\everycr={}}
\def\eqalignno#1{\displ@y \tabskip\plaincentering
  \halign to\displaywidth{\hfil$\@lign\displaystyle{##}$\tabskip\z@skip
    &$\@lign\displaystyle{{}##}$\hfil\tabskip\plaincentering
    &\llap{$\@lign##$}\tabskip\z@skip\crcr
    #1\crcr}}
\def\leqalignno#1{\displ@y \tabskip\plaincentering
  \halign to\displaywidth{\hfil$\@lign\displaystyle{##}$\tabskip\z@skip
    &$\@lign\displaystyle{{}##}$\hfil\tabskip\plaincentering
    &\kern-\displaywidth\rlap{$\@lign##$}\tabskip\displaywidth\crcr
    #1\crcr}}
\def\plainLet@{\relax\iffalse{\fi\let\\=\cr\iffalse}\fi}
\def\plainvspace@{\def\vspace##1{\noalign{\vskip##1}}}

\def\intic@{\mathchoice{\hskip5\p@}{\hskip4\p@}{\hskip4\p@}{\hskip4\p@}}
\def\negintic@
 {\mathchoice{\hskip-5\p@}{\hskip-4\p@}{\hskip-4\p@}{\hskip-4\p@}}
\def\intkern@{\mathchoice{\!\!\!}{\!\!}{\!\!}{\!\!}}
\def\intdots@{\mathchoice{\cdots}{{\cdotp}\mkern1.5mu
    {\cdotp}\mkern1.5mu{\cdotp}}{{\cdotp}\mkern1mu{\cdotp}\mkern1mu
      {\cdotp}}{{\cdotp}\mkern1mu{\cdotp}\mkern1mu{\cdotp}}}
\newcount\intno@
\def\iint{\intno@=\tw@\futurelet\next\ints@}
\def\iiint{\intno@=\thr@@\futurelet\next\ints@}
\def\iiiint{\intno@=4 \futurelet\next\ints@}
\def\idotsint{\intno@=\z@\futurelet\next\ints@}
\def\ints@{\findlimits@\ints@@}
\newif\iflimtoken@
\newif\iflimits@
\def\findlimits@{\limtoken@false\limits@false\ifx\next\limits
 \limtoken@true\limits@true\else\ifx\next\nolimits\limtoken@true\limits@false
    \fi\fi}
\def\multintlimits@{\intop\ifnum\intno@=\z@\intdots@
  \else\intkern@\fi
    \ifnum\intno@>\tw@\intop\intkern@\fi
     \ifnum\intno@>\thr@@\intop\intkern@\fi\intop}
\def\multint@{\int\ifnum\intno@=\z@\intdots@\else\intkern@\fi
   \ifnum\intno@>\tw@\int\intkern@\fi
    \ifnum\intno@>\thr@@\int\intkern@\fi\int}
\def\ints@@{\iflimtoken@\def\ints@@@{\iflimits@
   \negintic@\mathop{\intic@\multintlimits@}\limits\else
    \multint@\nolimits\fi\eat@}\else
     \def\ints@@@{\multint@\nolimits}\fi\ints@@@}
\def\Sb{_\bgroup\vspace@
        \baselineskip=\fontdimen10 \scriptfont\tw@
        \advance\baselineskip by \fontdimen12 \scriptfont\tw@
        \lineskip=\thr@@\fontdimen8 \scriptfont\thr@@
        \lineskiplimit=\thr@@\fontdimen8 \scriptfont\thr@@
        \Let@\vbox\bgroup\halign\bgroup \hfil$\scriptstyle
            {##}$\hfil\cr}
\def\endSb{\crcr\egroup\egroup\egroup}
\def\Sp{^\bgroup\vspace@
        \baselineskip=\fontdimen10 \scriptfont\tw@
        \advance\baselineskip by \fontdimen12 \scriptfont\tw@
        \lineskip=\thr@@\fontdimen8 \scriptfont\thr@@
        \lineskiplimit=\thr@@\fontdimen8 \scriptfont\thr@@
        \Let@\vbox\bgroup\halign\bgroup \hfil$\scriptstyle
            {##}$\hfil\cr}
\def\endSp{\crcr\egroup\egroup\egroup}
\def\Let@{\relax\iffalse{\fi\let\\=\cr\iffalse}\fi}
\def\vspace@{\def\vspace##1{\noalign{\vskip##1 }}}
\def\aligned{\,\vcenter\bgroup\plainvspace@\plainLet@\openup\jot\m@th\ialign
  \bgroup \strut\hfil$\displaystyle{##}$&$\displaystyle{{}##}$\hfil\crcr}
\def\endaligned{\crcr\egroup\egroup}
\def\matrix{\,\vcenter\bgroup\plainLet@\plainvspace@
    \normalbaselines
  \m@th\ialign\bgroup\hfil$##$\hfil&&\quad\hfil$##$\hfil\crcr
    \mathstrut\crcr\noalign{\kern-\baselineskip}}
\def\endmatrix{\crcr\mathstrut\crcr\noalign{\kern-\baselineskip}\egroup
                \egroup\,}
\newtoks\hashtoks@
\hashtoks@={#}
\def\format{\crcr\egroup\iffalse{\fi\ifnum`}=0 \fi\format@}
\def\format@#1\\{\def\preamble@{#1}%
  \def\c{\hfil$\the\hashtoks@$\hfil}%
  \def\r{\hfil$\the\hashtoks@$}%
  \def\l{$\the\hashtoks@$\hfil}%
  \setbox\z@=\hbox{\xdef\Preamble@{\preamble@}}\ifnum`{=0 \fi\iffalse}\fi
   \ialign\bgroup\span\Preamble@\crcr}

\def\cases{\left\{\,\vcenter\bgroup\plainvspace@
     \normalbaselines\openup\jot\m@th
      \plainLet@\ialign\bgroup$\displaystyle{##}$\hfil&\quad$\displaystyle{{}##}$\hfil\crcr
      \mathstrut\crcr\noalign{\kern-\baselineskip}}

\newif\iftagsleft@
\tagsleft@true
\def\TagsOnRight{\global\tagsleft@false}
\def\tag#1$${\iftagsleft@\leqno\else\eqno\fi
 \hbox{\def\pagebreak{\global\postdisplaypenalty-\@M}%
 \def\nopagebreak{\global\postdisplaypenalty\@M}\rm(#1\unskip)}%
  $$\postdisplaypenalty\z@\ignorespaces}
\interdisplaylinepenalty=\@M
\def\plainallowdisplaybreak@{\def\allowdisplaybreak{\noalign{\allowbreak}}}
\def\plaindisplaybreak@{\def\displaybreak{\noalign{\break}}}
\def\align#1\endalign{\def\tag{&}\plainvspace@\plainallowdisplaybreak@\plaindisplaybreak@
  \iftagsleft@\plainlalign@#1\endalign\else
   \plainralign@#1\endalign\fi}
\def\plainralign@#1\endalign{\displ@y\plainLet@\tabskip\plaincentering\halign to\displaywidth
     {\hfil$\displaystyle{##}$\tabskip=\z@&$\displaystyle{{}##}$\hfil
       \tabskip=\plaincentering&\llap{\hbox{\rm(##\unskip)}}\tabskip\z@\crcr
             #1\crcr}}
\def\plainlalign@
 #1\endalign{\displ@y\plainLet@\tabskip\plaincentering\halign to \displaywidth
   {\hfil$\displaystyle{##}$\tabskip=\z@&$\displaystyle{{}##}$\hfil
   \tabskip=\plaincentering&\kern-\displaywidth
        \rlap{\hbox{\rm(##\unskip)}}\tabskip=\displaywidth\crcr
               #1\crcr}}

\def\re@#1{\par\hangindent\parindent\indent\llap{#1\enspace}\ignorespaces}
\def\qfootnote#1{\edef\@sf{\spacefactor\the\spacefactor}{}#1\@sf
      \insert\footins{\let\egroup=}\footnotesize 
      \interlinepenalty100 \let\par=\endgraf
        \leftskip=0pt \rightskip=0pt
        \splittopskip=10pt plus 1pt minus 1pt \floatingpenalty=20000
   \smallskip\re@{#1}\bgroup\strut\aftergroup{\strut\egroup}\let\next}
\catcode`\@=\active
\TagsOnRight

\begin{document}
\title{\bf Universal bounds on eigenvalues of the buckling problem on spherical domains
\footnote{This research is supported by NSFC of China (No.
10671181), Project of Henan Provincial department of Sciences and
Technology (No. 092300410143), and NSF of Henan Provincial Education
department (No. 2009A110010).}}
\author{Guangyue Huang,\ Xingxiao Li
\footnote{The corresponding author. Email: xxl$@$henannu.edu.cn}\, , \  Linfen Cao\\
{\normalsize Department of Mathematics, Henan Normal University}
\\{\normalsize Xinxiang 453007, Henan, P.R. China}}
\date{}
\maketitle
\begin{quotation}
\noindent{\bf Abstract.} In this paper we study the eigenvalues of
buckling problem on domains in a unit sphere. By introducing a new
parameter and using Cauchy inequality, we optimize the
inequality obtained by Wang and Xia in \cite{wangxia07}.\\
{{\bf Keywords}: eigenvalue, universal bounds, buckling problem.} \\
{{\bf Mathematics Subject Classification}: Primary 35P15, Secondary
58G25.}

\end{quotation}

\section{Introduction}

Let $\Omega$ be a connected bounded domain in an
$n(\geq2)$-dimensional Euclidean space $\mathbb{R}^n$ and $\nu$ be
the unit outward normal vector field of $\partial\Omega$. The
well-known eigenvalue problem
\begin{equation}\label{Intro1}\left\{\begin{array}{ll}
\Delta^2 u=\Lambda(-\Delta) u & {\rm in}\ \Omega,\\
u=\frac{\partial u}{\partial\nu}=0 & {\rm on}\
\partial\Omega
\end{array}\right.
\end{equation}
is called a {\em buckling problem}, which is used to describe the
critical buckling load of a clamped plate subjected to a uniform
compressive force around its boundary.

Let $$0<\Lambda_1\leq\Lambda_2\leq \Lambda_3\leq \cdots$$ denote the
successive eigenvalues for \eqref{Intro1}, where each eigenvalue is
repeated according to its multiplicity. In 1956,
Payne-P\'{o}lya-Weinberger \cite{ppw2} proved
\begin{equation}\label{Intro2}
\Lambda_2\leq 3\Lambda_1\ \ \ {\rm for}\ \Omega\in \mathbb{R}^2.
\end{equation} Following the method of Payne-P\'{o}lya-Weinberger in \cite{ppw2},
it reads that for $\Omega\in\mathbb{R}^n$ (for the generalization of
\eqref{Intro2} to $n$ dimensions, see \cite{ash04}):
$$\Lambda_2\leq \left(1+\frac8{n+2}\right)\Lambda_1.
$$
Subsequently, in 1984, Hile and Yeh \cite{hile1} improved the above
inequality as follows:
$$
\Lambda_2\leq \frac{n^2+8n+20}{(n+2)^2}\Lambda_1.
$$
In 1998, Ashbaugh \cite{ash} obtained
$$\sum_{i=1}^n\Lambda_{i+1}\leq (n+4)\Lambda_1.$$
In a recent survey paper, answering a question of Ashbaugh in
\cite{ash}, Cheng-Yang \cite{chengyang06} proved the following
universal inequalities on eigenvalues for the eigenvalue problem
\eqref{Intro1}:
$$
\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\leq\frac{4(n+2)}{
n^2}\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)\Lambda_i. $$

It has become clear that many similar inequalities on eigenvalues of
Laplacian of Payne-P\'{o}lya-Weinberger rely on facts involving
operators, their commutators, and traces. For the related research
and improvement in this direction, see
\cite{harrell97,huang09,levitin,huangchen,harrell,wang07,cheng4} and
the references therein.

Let $x_1,x_2,\ldots,x_{n+1}$ be the standard Euclidean coordinate
functions of $\mathbb{R}^{n+1}$. Then the unit sphere is defined by
$$\mathbb{S}^n=\left\{ (x_1,x_2,\ldots, x_{n+1})\in\mathbb{R}^{n+1}\ ;\
\sum_{\alpha=1}^{n+1}x_\alpha^2=1\right\}.$$ In 2007, Wang and Xia
\cite{wangxia07} considered the buckling problem on domains in a
unit sphere $\mathbb{S}^n$ and obtained the following result:

\begin{theorem}{\rm\cite{wangxia07}} Let $\Lambda_i$ be the
$i^{th}$ eigenvalue of the following eigenvalue problem:
$$\Delta^2 u=\Lambda(-\Delta) u \ \ {\rm in}\ \Omega,\ \ \ \
u=\frac{\partial u}{\partial\nu}=0\ \ {\rm on}\
\partial\Omega,$$ where $\Omega$ is a connected domain in
a unit sphere $\mathbb{S}^n(n\geq2)$ with smooth boundary
$\partial\Omega$ and $\nu$ is the unit outward normal vector field
of $\partial\Omega$. Then for any $\delta>0$,
\begin{equation}\label{Intro3}
\aligned 2\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\leq&
\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\left(\delta\Lambda_i
+\frac{\delta^2(\Lambda_i-(n-2))}{4(\delta\Lambda_i+n-2)}\right)\\
&+
\frac{1}{\delta}\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)\left(\Lambda_i
+\frac{(n-2)^2}{4}\right).
\endaligned
\end{equation}
\end{theorem}

We remark that the right hand side of inequality \eqref{Intro3}
depends on $\delta$. In the current paper, by introducing a new
parameter and using Cauchy inequality, we obtain a stronger
inequality than \eqref{Intro3} which is independent of $\delta$, and
derive an inequality of the type of Yang (see inequality
\eqref{Intro5}). Our main results are stated as follows:

\begin{theorem}\label{th:main} Let $\Lambda_i$ be the $i^{th}$ eigenvalue
of the following eigenvalue problem:
$$\Delta^2 u=\Lambda(-\Delta) u \ \ {\rm in}\ \Omega,\ \ \ \
u=\frac{\partial u}{\partial\nu}=0\ \ {\rm on}\
\partial\Omega,$$ where $\Omega$ is a connected domain in
a unit sphere $\mathbb{S}^n(n\geq2)$ with smooth boundary
$\partial\Omega$ and $\nu$ is the unit outward normal vector field
of $\partial\Omega$. Then
\begin{alignat}{1}\label{Intro4}
&\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2
\left(2+\frac{n-2}{\Lambda_i-(n-2)}\right)\\
\leq&
2\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\left(\Lambda_i
-\frac{n-2}{\Lambda_i-(n-2)}\right)\right\}^{1/2}\\
&\qquad \qquad\qquad
\times\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)
\left(\Lambda_i+\frac{(n-2)^2}{4}\right)\right\}^{1/2}.
\end{alignat}
\end{theorem}

\begin{corollary} Under the assumptions of Theorem 1.2,
\begin{equation}\label{Intro5}
\aligned\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\leq
\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)\left(\Lambda_i
-\frac{n-2}{\Lambda_i-(n-2)}\right)
\left(\Lambda_i+\frac{(n-2)^2}{4}\right),
\endaligned
\end{equation}
\begin{equation}\label{Intro6}\Lambda_{k+1}\leq
S_{k+1}+\sqrt{S_{k+1}^2-T_{k+1}},
\end{equation}
\begin{equation}\label{Intro7}\Lambda_{k+1}-\Lambda_k\leq2\sqrt{S_{k+1}^2-T_{k+1}},
\end{equation}
where \begin{equation}\label{Intro8}S_{k+1}=\frac1{
k}\sum_{i=1}^k\Lambda_i+\frac{1}{
2k}\sum_{i=1}^k\left(\Lambda_i-\frac{n-2}{\Lambda_i-(n-2)}\right)
\left(\Lambda_i+\frac{(n-2)^2}{4}\right),\end{equation}
\begin{equation}\label{Intro9}T_{k+1}=\frac{1}{k}\sum_{i=1}^k\Lambda_i^2+\frac{1}{
k}\sum_{i=1}^k\Lambda_i\left(\Lambda_i-\frac{n-2}{\Lambda_i-(n-2)}\right)
\left(\Lambda_i+\frac{(n-2)^2}{4}\right).\end{equation}\end{corollary}

\begin{remark} {\rm The inequality \eqref{Intro4} is equivalent to
$$\aligned 2\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2
\leq&-\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\frac{n-2}{\Lambda_i-(n-2)}\\
&+
2\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2
\left(\Lambda_i-\frac{n-2}{\Lambda_i-(n-2)}\right)\right\}^{1/2}\\
&\times\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)
\left(\Lambda_i+\frac{(n-2)^2}{4}\right)\right\}^{1/2}.\endaligned$$
From the inequality \eqref{add} in Section 2 and making use of
Cauchy inequality, the following inequality can be deduced:
$$\aligned
-\sum_{i=1}^k&(\Lambda_{k+1}-\Lambda_i)^2\frac{n-2}{\Lambda_i-(n-2)}\\
&+
2\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\left(\Lambda_i
-\frac{n-2}{\Lambda_i-(n-2)}\right)\right\}^{1/2}\\
&\times\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)
\left(\Lambda_i+\frac{(n-2)^2}{4}\right)\right\}^{1/2}\\
\leq&-\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\frac{n-2}{\Lambda_i-(n-2)}\\
&+
\delta\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\left(\Lambda_i
-\frac{n-2}{\Lambda_i-(n-2)}\right)\\
&+\frac{1}{\delta}\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)
\left(\Lambda_i+\frac{(n-2)^2}{4}\right)\\
=&\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\left(\delta\Lambda_i
-\frac{(\delta+1)(n-2)}{\Lambda_i-(n-2)}\right)\\
&+
\frac{1}{\delta}\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)
\left(\Lambda_i+\frac{(n-2)^2}{4}\right)\\
\leq&\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\left(\delta\Lambda_i
+\frac{\delta^2(\Lambda_i-(n-2))}{4(\delta\Lambda_i+n-2)}\right)\\
&+ \frac{1}{\delta}\sum_{i=1}^k(\Lambda_{k+1}
-\Lambda_i)\left(\Lambda_i+\frac{(n-2)^2}{4}\right),
\endaligned$$
which shows that the inequality \eqref{Intro4} is sharper than
inequality \eqref{Intro3}. Therefore, Theorem 1.2 improves Theorem
1.1.}\end{remark}

\section{Proof of Theorem}

By the method of constructing trial functions introduced by Cheng
and Yang in \cite{chengyang06}, for any $\alpha=1,\ldots, n+1$ and
each $i=1,\ldots,k$, the vector-valued functions $x_\alpha\nabla
u_i$ can be decomposed as
\begin{equation}\label{2sec1}x_\alpha\nabla u_i=\nabla h_{\alpha i}+W_{\alpha i},\end{equation}
where $h_{\alpha i}\in H^2_{2,D}(\Omega)$, $\nabla h_{\alpha i}$ is
the projection of $x_\alpha\nabla u_i$ in
$\mathbf{H}^2_{1,D}(\Omega)$ and $W_{\alpha i}\perp
\mathbf{H}^2_{1,D}(\Omega)$ (for the definitions of
$H^2_{2,D}(\Omega)$ and $\mathbf{H}^2_{1,D}(\Omega)$, we refer to
\cite{wangxia07}). Hence,
$$W_{\alpha i}|_{\partial\Omega}=0$$ and
$$ \int_\Omega\langle W_{\alpha i},\nabla u\rangle=0, \ \ \ \ {\rm for\ any}\ u\in  H^2_{2,D}(\Omega).$$

Define $\|f\|^2=\int_\Omega|f|^2$. Then (see inequalities (2.19) and
(2.40) in \cite{wangxia07})
\begin{equation}\label{2sec2}
\|x_\alpha\nabla u_i\|^2=\|\nabla h_{\alpha i}\|^2+\|W_{\alpha
i}\|^2
\end{equation}
and for any $\delta>0$,
\begin{alignat}{1}\label{2sec3}
\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2r_{\alpha
i}\leq&\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\Big(\delta p_{\alpha
i}+(\delta\Lambda_i+n-2)\|W_{\alpha i}\|^2\\
&+\delta\|\langle\nabla x_\alpha,\nabla u_i\rangle\|^2 \Big)
+\frac{1}{\delta}\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)\|Z_{\alpha
i}\|^2,\end{alignat}
where
$$\aligned
r_{\alpha i}=&2\|\langle\nabla x_\alpha,\nabla
u_i\rangle\|^2+\int_\Omega \langle\nabla x_\alpha^2,\Delta u_i\nabla
u_i\rangle+(n-2)\| x_\alpha\nabla u_i\|^2,\\p_{\alpha
i}=&\int_\Omega\langle\nabla x_\alpha^2,u_i\nabla (\Delta
u_i)+\Lambda_iu_i\nabla u_i\rangle,\\
Z_{\alpha i}=&\nabla\langle\nabla x_\alpha,\nabla
u_i\rangle-\frac{n-2}{ 2}x_\alpha\nabla u_i.\endaligned
$$

Next, we are to optimize the inequality (2.41) in \cite{wangxia07}.
Let $C$ be a positive constant. Then it follows from Cauchy
inequality that
\begin{alignat}{1}\label{2sec4} C\|\langle\nabla x_\alpha,\nabla
u_i\rangle\|^2=&C\int_\Omega\langle\nabla x_\alpha,\nabla
u_i\rangle^2\\
=&-C\int_\Omega  x_\alpha {\rm div}(\langle\nabla
x_\alpha,\nabla u_i\rangle\nabla u_i)\\
=&-C\int_\Omega\langle x_\alpha\nabla u_i,\nabla \langle\nabla
x_\alpha,\nabla u_i\rangle\rangle-C\int_\Omega\langle\nabla
x_\alpha,\nabla u_i\rangle x_\alpha\Delta u_i\\
=&-C\int_\Omega\langle \nabla h_{\alpha i},\nabla \langle\nabla
x_\alpha,\nabla
u_i\rangle\rangle-\frac{C}{2}\int_\Omega\langle\nabla
x_\alpha^2,\nabla u_i\rangle \Delta u_i\\
\leq&(\delta\Lambda_i+n-2)\|\nabla h_{\alpha
i}\|^2+\frac{C^2}{4(\delta\Lambda_i+n-2)}\|\nabla \langle\nabla
x_\alpha,\nabla u_i\rangle\|^2\\
&-\frac{C}{2}\int_\Omega\langle\nabla x_\alpha^2,\nabla u_i\rangle
\Delta u_i,
\end{alignat}
where ${\rm div}(Z)$ denotes the divergence of $Z$. Applying
\eqref{2sec4} to \eqref{2sec3} yields
\begin{alignat}{1}\label{2sec5}
\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2r_{\alpha
i}\leq&\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\Big(\delta p_{\alpha
i}+(\delta\Lambda_i+n-2)\|W_{\alpha
i}\|^2\\&+(\delta-C)\|\langle\nabla x_\alpha,\nabla
u_i\rangle\|^2+C\|\langle\nabla x_\alpha,\nabla
u_i\rangle\|^2\Big)\\
&+\frac{1}{\delta}\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)\|Z_{\alpha
i}\|^2\\
\leq&\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\Big(\delta p_{\alpha
i}+(\delta\Lambda_i+n-2)(\|W_{\alpha i}\|^2+\|\nabla h_{\alpha
i}\|^2)\\&+(\delta-C)\|\langle\nabla x_\alpha,\nabla
u_i\rangle\|^2+\frac{C^2}{4(\delta\Lambda_i+n-2)}\|\nabla
\langle\nabla x_\alpha,\nabla
u_i\rangle\|^2\\
&-\frac{C}{2}\int_\Omega\langle\nabla x_\alpha^2,\nabla u_i\rangle
\Delta
u_i\Big)+\frac{1}{\delta}\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)\|Z_{\alpha
i}\|^2\\
=&\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\Big(\delta p_{\alpha
i}+(\delta\Lambda_i+n-2)\|x_\alpha\nabla
u_i\|^2\\&+(\delta-C)\|\langle\nabla x_\alpha,\nabla
u_i\rangle\|^2+\frac{C^2}{4(\delta\Lambda_i+n-2)}\|\nabla
\langle\nabla x_\alpha,\nabla
u_i\rangle\|^2\\
&-\frac{C}{2}\int_\Omega\langle\nabla x_\alpha^2,\nabla u_i\rangle
\Delta
u_i\Big)+\frac{1}{\delta}\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)\|Z_{\alpha
i}\|^2,
\end{alignat}
where in the last step in \eqref{2sec5} we have used equality
\eqref{2sec2}. A direct calculation yields (see (2.44), (2.45),
(2.46) and (2.47) in \cite{wangxia07})
\begin{equation}\label{2sec6}
\sum_{\alpha=1}^{n+1}r_{\alpha i}=n,
\end{equation}
\begin{equation}\label{2sec7}
\sum_{\alpha=1}^{n+1}p_{\alpha i}=0,
\end{equation}
\begin{equation}\label{2sec8}
\sum_{\alpha=1}^{n+1}\|x_\alpha\nabla
u_i\|^2=\sum_{\alpha=1}^{n+1}\|\langle\nabla x_\alpha,\nabla
u_i\rangle\|^2=1,
\end{equation}
\begin{equation}\label{2sec9}
\sum_{\alpha=1}^{n+1}\|\nabla \langle\nabla x_\alpha,\nabla
u_i\rangle\|^2=\Lambda_i-(n-2)
\end{equation}
and
\begin{equation}\label{2sec10}
\sum_{\alpha=1}^{n+1}\|Z_{\alpha i}\|^2=\Lambda_i+\frac{(n-2)^2}{
4}.
\end{equation}
Therefore, summing up \eqref{2sec5} over $\alpha$ from 1 to $n+1$,
one gets
$$\aligned
n\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\leq&\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2
\Big(\delta\Lambda_i+n-2+(\delta-C)\\
&+\frac{C^2}{4(\delta\Lambda_i+n-2)}(\Lambda_i-(n-2))\Big)\\
&+\frac1\delta\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)\left(\Lambda_i+\frac{(n-2)^2}{
4}\right).
\endaligned$$
That is,
\begin{alignat}{1}\label{2sec11}
2\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\leq&\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2
\left(\delta\Lambda_i+(\delta-C)\right.\\
&\left.+\frac{C^2}{
4(\delta\Lambda_i+n-2)}(\Lambda_i-(n-2))\right)\\
&
+\frac{1}{\delta}\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)\left(\Lambda_i+\frac{(n-2)^2}{
4}\right).
\end{alignat}
Taking $$C=\frac{2(\delta\Lambda_i+n-2)}{\Lambda_i-(n-2)}$$ in
\eqref{2sec11} yields
\begin{alignat}{1}\label{2sec12}
2\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\leq&\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2
\left(\delta\Lambda_i-\frac{(\delta+1)(n-2)}{\Lambda_i-(n-2)}\right)\\
&+\frac{1}{\delta}\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)\left(\Lambda_i+\frac{(n-2)^2}{
4}\right),
\end{alignat}
and hence
\begin{alignat}{1}\label{2sec13}
&\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2
\left(2+\frac{n-2}{\Lambda_i-(n-2)}\right)\\
\leq& \delta\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\left(\Lambda_i
-\frac{n-2}{\Lambda_i-(n-2)}\right)\\
&+\frac{1}{\delta}\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)
\left(\Lambda_i+\frac{(n-2)^2}{ 4}\right).
\end{alignat}

To complete the proof of Theorem \ref{th:main}, we need the
following lemma:

\begin{lemma} Let $\Omega$ be a connected bounded domain in
$\mathbb{S}^n$. Then
\begin{equation}\label{2sec14}\Lambda_1\geq
n.\end{equation}\end{lemma}

\begin{proof} Let $\Omega_1$, $\Omega_2$ be two connected bounded
domains in $\mathbb{S}^n$ and $\Omega_1\subset\Omega_2$. Let
$u_1(\Omega_1)$ be the eigenfunction corresponding to
$\Lambda_1(\Omega_1)$. Then the function defined by $$\widetilde{
u_1}=\left\{\begin{array}{ll} u_1 &\ \ \  {\rm in }\ \ \Omega_1,\\
0 &\ \ \  {\rm in }\ \ \Omega_2- \Omega_1
\end{array}\right.
$$
is a eigenfunction corresponding to $\Lambda_1(\Omega_2)$. Hence,
$\Lambda_1(\Omega_1)\geq \Lambda_1(\Omega_2)$. Denote by $\lambda_1$
the first eigenvalue of Laplacian. It is easy to see
$\Lambda_1(\mathbb{S}^n)=\lambda_1(\mathbb{S}^n)=n$ because there
are no boundary conditions in this case. It follows that
$\Lambda_1(\Omega)\geq \Lambda_1(\mathbb{S}^n)=n$ by setting
$\Omega=\Omega_1$ and $\Omega_2=\mathbb{S}^n$. This completes the
proof of Lemma 2.1.\end{proof}

Inequality \eqref{2sec14} shows that for any $i$,
\begin{equation}\label{add}\Lambda_i-\frac{n-2}{\Lambda_i-(n-2)}>0.\end{equation}
Minimizing the right hand side of \eqref{2sec13} as a function of
$\delta$ by choosing
$$\delta=\left(\frac{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)
\left(\Lambda_i+\frac{(n-2)^2}{
4}\right)}{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2
\left(\Lambda_i-\frac{n-2}{\Lambda_i-(n-2)}\right)}\right)^{1/2},$$
we obtain the inequality \eqref{Intro4}, completing the proof of
Theorem \ref{th:main}.

\vspace*{3mm}

{\bf Proof of Corollary 1.3.} It is easy to see from \eqref{Intro4}
that
\begin{alignat}{1}\label{2sec15}
\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\leq&
\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2
\left(\Lambda_i-\frac{n-2}{\Lambda_i-(n-2)}\right)\right\}^{1/2}\\
&\qquad\qquad\times\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)
\left(\Lambda_i+\frac{(n-2)^2}{4}\right)\right\}^{1/2}.
\end{alignat}
One can check by induction that
$$\aligned
&\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2
\left(\Lambda_i-\frac{n-2}{\Lambda_i-(n-2)}\right)\right\}
\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)
\left(\Lambda_i+\frac{(n-2)^2}{
4}\right)\right\}\\
\leq&\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)^2\right\}
\left\{\sum_{i=1}^k(\Lambda_{k+1}-\Lambda_i)
\left(\Lambda_i-\frac{n-2}{\Lambda_i-(n-2)}\right)\left(\Lambda_i+\frac{(n-2)^2}{
4}\right)\right\},
\endaligned$$
which together with inequality \eqref{2sec15} yields inequality
\eqref{Intro5}.

Solving the quadratic polynomial of $\Lambda_{k+1}$ in
\eqref{Intro5}, we obtain inequality \eqref{Intro6}.

Replacing $k+1$ with $k$ in \eqref{Intro5}, we obtain
$$\sum_{i=1}^{k-1}(\Lambda_k-\Lambda_i)^2\leq
\sum_{i=1}^{k-1}(\Lambda_k-\Lambda_i)\left(\Lambda_i-\frac{n-2}{\Lambda_i-(n-2)}\right)
\left(\Lambda_i+\frac{(n-2)^2}{4}\right).$$ Therefore,
$$\sum_{i=1}^k(\Lambda_k-\Lambda_i)^2\leq
\sum_{i=1}^k(\Lambda_k-\Lambda_i)\left(\Lambda_i-\frac{n-2}{\Lambda_i-(n-2)}\right)
\left(\Lambda_i+\frac{(n-2)^2}{4}\right).$$ Solving this inequality
of quadratic polynomial for $\Lambda_{k}$, we infer
\begin{equation}\label{2sec16}\Lambda_k\geq
S_{k+1}-\sqrt{S_{k+1}^2-T_{k+1}},\end{equation} where $S_{k+1},
T_{k+1}$ are given by \eqref{Intro8} and \eqref{Intro9}
respectively. Therefore, the inequality \eqref{Intro7} follows from
\eqref{2sec16} and \eqref{Intro6}. Then Corollary 1.3 is completed.

\end{document}